# On Supports for Graphs of Bounded Genus [^1] A preliminary version of these results appeared in EUROCOMB’23 [raman2023hypergraph]

### v3

@@ 1 Introduction
A hypergraph $(X,\mathcal{E})$ consists of a set $X$ and a collection $\mathcal{E}$ of subsets of $X$ called hyperedges. Hypergraphs generalize graphs (where $\mathcal{E}$ consists only of 2-element subsets of $X$). We are interested in questions of representing hypergraphs by sparse graphs that capture some of its structure. The notion of structure we want to preserve is that of connectedness of the hyperedges.
Consider a hypergraph $(X,\mathcal{E})$. A graph $Q$ on $X$ is a support for the hypergraph if for each $E\in\mathcal{E}$, the induced graph $Q[E]$ of $G$ is connected. If we forgo sparsity, we can trivially construct a support by taking a complete graph on $X$. The problems become non-trivial if we impose sparsity. For example, if we require the support graph to be planar, then even deciding if a hypergraph admits such a support is NP-hard [Johnson1987Hypergraph].
The notion of a support was introduced by Voloshina and Feinberg [voloshina1984planarity] in the context of VLSI circuits, and has since been rediscovered by several communities. For example, the existence and construction of sparse supports arises naturally in the problem of visualizing hypergraphs [bereg2015colored, bereg2011red, brandes2010blocks, brandes2012path, buchin2011planar, havet2022overlaying, hurtado2018colored], and in the design of networks [chen2015polynomial, hosoda2012approximability, korach2003clustering, onus2011minimum].
Since it is already NP-hard to decide if a hypergraph admits a support that is a planar graph, it is natural to consider restricted settings that ensure the existence of a sparse support. Raman and Ray [RR18] considered hypergraphs defined by non-piercing regions [^3] A slightly imprecise definition is the following: a family $\mathcal{R}$ of path connected regions in the plane or on an oriented surface such that for any two regions $R,R^{\prime}\in\mathcal{R}$, the regions $R\setminus R^{\prime}$ and $R^{\prime}\setminus R$ are also connected are called non-piercing regions. They include disks, pseudodisks, unit height-axis parallel rectangles, etc. See Def. 17. in the plane (see Section 7.2 for a definition of these hypergraphs). The authors showed the existence of planar primal, dual and intersection supports for these hypergraphs.
@@
Figure 1: Support for hypergraph defined by disks and points in the plane.
@@
Our results on the existence of supports of bounded genus implies, via the local-search framework, PTAS for the problems studied by [RR18] generalized to surfaces of bounded genus. We also show that the non-piercing condition is not sufficient to guarantee PTAS for hypergraphs on surfaces of bounded genus by showing an APX-hardness result for a covering problem defined by non-piercing regions on a torus.
@@ 1.1 Problems studied
We now define intersection hypergraphs that are a common generalization of primal and dual hypergraphs. Let $G=(V,E)$ be a graph and let $\mathcal{H}$ and $\mathcal{K}$ be two families of connected subgraphs of $G$. An intersection hypergraph is the hypergraph $(\mathcal{H},\{\mathcal{H}_{K}\}_{K\in\mathcal{K}})$, i.e., a hypergraph on $\mathcal{H}$ such that each $K\in\mathcal{K}$ defines a hyperedge consisting of the subgraphs in $\mathcal{H}$ intersecting at a vertex of $K$. As in the dual setting, choosing a subset of $\mathcal{H}$ as terminals does not add any additional complication, and hence we define the intersection hypergraph on all of $\mathcal{H}$. Setting either $\mathcal{H}$ or $\mathcal{K}$ as singleton vertices yields respectively, the primal and dual hypergraphs. An intersection support is a support for the intersection hypergraph.
@@
Our main result is the following: We show that if the host graph has genus $g$ and the subgraphs in $\mathcal{H}$ and $\mathcal{K}$ are cross-free, then there is an intersection support of genus $g$. This implies an equivalent result for the primal and dual hypergraphs. Figure 2 shows examples of primal, dual and intersection supports.
@@
Figure 2: (a) and (b): Primal and Dual hypergraphs on the graph system $(G,\mathcal{H})$ where $G$ is the graph on vertices $\{a,b\ldots,f\}$, and $\mathcal{H}=\{H_{1},\ldots H_{4}\}$ with $H_{1}=\{a,b,c,d\}$, $H_{2}=\{c,d,e\}$, $H_{3}=\{a,b,f,e\}$, $H_{4}=\{a,b,c,e\}$; (c): Intersection hypergraph for $(G,\mathcal{H}^{\prime},\mathcal{K})$ with $\mathcal{H}^{\prime}=\{H^{\prime}_{1},\ldots,H^{\prime}_{4}\}$ and $\mathcal{K}=\{K_{1},\ldots,K_{3}\}$ where $H^{\prime}_{1}=\{a,b,c\}$, $H^{\prime}_{2}=\{a,b,c,e\}$, $H^{\prime}_{3}=\{c,e,f\}$, $H^{\prime}_{4}=\{b,e,d\}$, and $K_{1}=\{c,d\}$, $K_{2}=\{b,f\}$, $K_{3}=\{e,d\}$. Figures in (d)-(f) show the respective support graphs.
@@ 1.2 Related Work
The notion of a planar analogue of a hypergraph was first suggested by Zykov [zykov1974hypergraphs], who defined a hypergraph to be planar if there is a plane graph on the vertices of the hypergraph such that for each hyperedge, there is a bounded face of the embedding containing only the elements of this hyperedge. Equivalently, a hypergraph is Zykov-planar iff its incidence bipartite graph is planar. Voloshina and Feinberg [voloshina1984planarity] introduced the notion of hypergraph planarity in the context of planarizing VLSI circuits (see the monograph by Feinberg et al., [feinberg2012vlsi] and references therein). The notion is now called a planar support. Johnson and Pollak [Johnson1987Hypergraph] showed that the problem of deciding if a hypergraph has a planar support is NP-hard.
@@
Since then, several authors have studied the question of deciding if a hypergraph admits a support from a restricted family of graphs. Tarjan and Yannakakis [tarjan1984simple] showed that we can decide in linear time if a hypergraph admits a tree support. Buchin et al., [buchin2011planar] showed polynomial time algorithms to decide if a hypergraph admits a support that is a path, a cycle, or a tree with bounded maximum degree. Further, the authors sharpen the result of Johnson and Pollak [Johnson1987Hypergraph] by showing that deciding if a hypergraph admits a support that is a 2-outerplanar [^4] A graph is 2-outerplanar if the graph can be embedded in the plane such that the vertices are on two concentric circles and removing all vertices of outer face results in an outerplanar graph. graph is NP-hard. The problems of constructing a support with the fewest number of edges, or with minimum maximum degree have also been studied in [baldoni2007tera, onus2011minimum].
@@
As stated earlier, the existence of supports and related notions have been used in the analysis of algorithms under a local-search framework for packing and covering problems on geometric hypergraphs. Chan and Har-Peled [ChanH12] gave a PTAS for the Independent Set problem for an arrangement of pseudodisks in the plane. Mustafa and Ray [mustafa2010improved] gave a PTAS for the Hitting Set problem for a set of points and pseudodisks in the plane (both results apply more generally to k-admissible regions  [^5] A set of connected bounded regions $\mathcal{R}$ in the plane, each of whose boundary is a simple Jordan curve is $k$-admissible (for even $k$) if for any $^{\prime}\in\mathcal{R}, RRR\setminus R^{\prime}$ and $R^{\prime}\setminus R$ are connected, their boundaries are in general position, and intersect each other at most $k$ times. If $k=2$, the regions are called pseudodisks.[whitesides1990k] ). The results of [ChanH12, mustafa2010improved] were extended by Basu-Roy et al., [BasuRoy2018] who used the same framework to design a PTAS for the Set Cover, and Dominating Set problem for the intersection graph of pseudodisks. Raman and Ray [RR18] showed the existence of a planar support for intersection hypergraphs of non-piercing regions (see Section 7 for the definition), and their result gave a unified analysis for all packing and covering problems described above. The local-search framework has also found applications in variants of the art gallery/terrain guarding problem [krohn2014guarding, DBLP:conf/wads/BandyapadhyayR17], where the geometric objects are defined implicitly as the visibility region of a guard, and the regions to be guarded are continuous regions.
@@
Cabello and Gajser [CabelloG14] showed that Independent Set and Vertex Cover admit a PTAS on graphs excluding a fixed minor, and Aschner et al., [DBLP:conf/walcom/AschnerKMY13] studied some packing and covering problems involving geometric non-planar graphs. Unlike their results, for the problems we study, we require additional work to show the existence of an appropriate local-search graph and this, in some cases requires the construction of a support graph.
Besides packing and covering problems, we also consider coloring problems on hypergraphs. Keller and Smorodinsky [KellerS18] considered intersection hypergraphs of two families of disks, and showed that such hypergraphs admit a conflict-free coloring [^6] In a conflict-free coloring, we want to color the vertices of the hypergraph such that each hyperedge has a uniqely colored vertex. with $O(\log n)$ colors. Keszegh [Keszegh20] generalized their result to intersection hypergraphs of two families of pseudodisks and showed that these hypergraphs admit a conflict-free coloring with $O(\log n)$ colors. The results for conflict-free coloring follow by showing that the hypergraph can be colored with a constant number of colors, i.e., a coloring where no hyperedge is monochromatic.
@@ 2 Preliminaries
For a vertex $v\in V(G)$, we use $N_{G}(v)$ to denote the neighbors of $v$ in $G$ (or just $N(v)$ if $G$ is clear from context). We use $u\sim v$ or $e\sim v$ to denote respectively, if $u$ and $v$ are adjacent in $G$, or an edge $e$ incident to vertex $v$. For $S\subseteq V$, we use $G[S]$ to denote the subgraph of $G$ induced on $S$. We use $\mathcal{H}_{v}=\{H\in\mathcal{H}:v\in H\}$. Similarly, let $\mathcal{H}_{e}=\{H\in\mathcal{H}:e\in H\}$. We let $0pt(v)=|\{H\in\mathcal{H}:v\in H\}|$. Similarly, let $0pt(e)=|\{H\in\mathcal{H}:e\in H|$. We also use the notations $H\cap K$, $H\setminus K$, and $H\subseteq K$ to mean $V(H)\cap V(K)$, $V(H)\setminus V(K)$, and $V(H)\subseteq V(K)$ respectively for any two subgraphs $H,K$ of a graph $G$.
@@ 3.2 Algorithms for Packing and Covering Problems
For a graph system $(G,\mathcal{H})$, an intersection graph is the graph $D=(\mathcal{H},E)$, where $\{H,H^{\prime}\}\in E(D)$ if and only $V(H)$ and $V(H^{\prime})$ intersect at a vertex of $G$.
@@ 4 Non-piercing and Cross-free Systems
Let $\mathcal{R}$ be an arrangement of connected and bounded regions in the plane, and $\partial(R)$ defines the boundary of $R\in\mathcal{R}$. Then $\mathcal{R}\setminus\bigcup_{R\in\mathcal{R}}\partial(R)$ splits the plane into set $\mathcal{C}$ of connected components, each called a cell in the arrangement $\mathcal{R}$. Then the dual arrangement graph $G=(\mathcal{C},E)$ of $\mathcal{R}$ is a plane graph with vertex set $\mathcal{C}$ where two vertices are adjacent if and only of their corresponding cells are separated by boundary of some region in $\mathcal{R}$.
@@ 5 Vertex Bypassing
Let $(G,\mathcal{H})$ be a cross-free system with a cross-free embedding of $G$ with respect to $\mathcal{H}$. Suppose we apply VB(v)(v) to vertex $v\in V(G)$. Then, each subgraph $H$ in $(G^{\prime},\mathcal{H}^{\prime})$ is connected. Further, VB(v)(v) can be done in time $O(|\mathcal{H}||V(G)|^{4})$.
@@
Consider an $x^{\prime}\in V(G^{\prime})$ and its corresponding vertex in $R_{G^{\prime}}(H^{\prime}_{1},H^{\prime}_{2})$. First, suppose $v\not\in V(H_{1})\cap V(H_{2})$. If $x^{\prime}\not\in\{u_{0},\ldots,u_{k-1}\}$, then the cyclic-pattern at $\tilde{x^{\prime}}$ in $R_{G^{\prime}}(H^{\prime}_{1},H^{\prime}_{2})$ is isomorphic to the cyclic-pattern at $\tilde{x}$ (replacing $H^{\prime}_{j}$ by $H_{j}$, $j=1,2$), where $x$ the vertex in $G$ corresponding to $x^{\prime}$. Since $H_{1}$ and $H_{2}$ are cross-free at $x$, so are $H^{\prime}_{1}$ and $H^{\prime}_{2}$ at $x^{\prime}$. Since the subgraphs at $u_{i}$ are a subset of the subgraphs at $v$, it implies that no $u_{i}$ is contained in both $H_{1}$ and $H_{2}$. Hence, $H_{1}$ and $H_{2}$ are cross-free at each $u_{i}$.
Now, if $v\in V(H_{1})\cap V(H_{2})$, then for any vertex $y^{\prime}$ in $G^{\prime}$ so that its copy $y$ in $G$ is not connected to $v$ via a path that lies in $V(H_{1})\cap V(H_{2})$, the cyclic-pattern at $\tilde{y^{\prime}}$ and $\tilde{y}$ are isomorphic. If $y$ lies on path in $V(H_{1})\cap V(H_{2})$ to $v$, then $\tilde{y}=\tilde{v}$ in $R_{G}(H_{1},H_{2})$. In this case, observe that the cyclic pattern at $\tilde{y^{\prime}}$ is a sub-sequence of the cyclic pattern at $\tilde{y}$ since $\tilde{y^{\prime}}=\tilde{u_{i}}$ for some $u_{i}\in H^{\prime}_{1}\cap H^{\prime}_{2}$. Since $H_{1}$ and $H_{2}$ are non-crossing at $v$, it implies that $H^{\prime}_{1}$ and $H^{\prime}_{2}$ are non-crossing at $y^{\prime}$.
Since $H^{\prime}_{1}$ and $H^{\prime}_{2}$ are non-crossing at all vertices of $G^{\prime}$, and $H_{1},H_{2}\in\mathcal{H}$ were arbitrary, it implies $(G^{\prime},\mathcal{H}^{\prime})$ is cross-free.

### v5

@@ 1 Introduction
A hypergraph $(X,\mathcal{E})$ consists of a set $X$ and a collection $\mathcal{E}$ of subsets of $X$ called hyperedges. Hypergraphs generalize graphs (where $\mathcal{E}$ consists only of 2-element subsets of $X$). We are interested in questions of representing hypergraphs by sparse graphs that capture some of their structure. The notion of structure we want to preserve is that of connectedness of the hyperedges.
Consider a hypergraph $(X,\mathcal{E})$. A graph $Q$ on $X$ is a support for the hypergraph if for each $E\in\mathcal{E}$, the induced graph $Q[E]$ of $Q$ is connected. If we ignore the sparsity, we can trivially construct a support by taking a complete graph on $X$. The problems become non-trivial if we impose sparsity. For example, if we require the support graph to be planar, then even deciding if a hypergraph admits such a support is NP-hard [Johnson1987Hypergraph].
The notion of a support was introduced by Voloshina and Feinberg [voloshina1984planarity] in the context of VLSI circuits, and has since been rediscovered by several communities. For example, the existence and construction of sparse supports arise naturally in the problem of visualizing hypergraphs [bereg2015colored, bereg2011red, brandes2010blocks, brandes2012path, buchin2011planar, havet2022overlaying, hurtado2018colored], and in the design of networks [chen2015polynomial, hosoda2012approximability, korach2003clustering, onus2011minimum].
Since it is already NP-hard to decide if a hypergraph admits a support that is a planar graph, it is natural to consider restricted settings that ensure the existence of a sparse support. Raman and Ray [RR18] considered hypergraphs defined by non-piercing regions [^3] A slightly imprecise definition is the following: a family $\mathcal{R}$ of path-connected regions in the plane or on an oriented surface such that for any two regions $R,R^{\prime}\in\mathcal{R}$, the regions $R\setminus R^{\prime}$ and $R^{\prime}\setminus R$ are also connected, is called a family of non-piercing regions. They include disks, pseudodisks, unit height-axis parallel rectangles, etc. See Def. 17. in the plane (see Section 7.2 for a definition of these hypergraphs). The authors showed the existence of planar primal, dual and intersection supports for these hypergraphs.
@@
Figure 1: Support for a hypergraph defined by disks and points in the plane.
@@
Our results on the existence of supports of bounded genus imply, via the local-search framework, PTAS for the problems studied by [RR18] generalized to surfaces of bounded genus. We also show that the non-piercing condition is not sufficient to guarantee PTAS for hypergraphs on surfaces of bounded genus by showing an APX-hardness result for a covering problem defined by non-piercing regions on a torus.
@@ 1.1 Problems studied
We now define intersection hypergraphs that are a common generalization of primal and dual hypergraphs. Let $G=(V,E)$ be a graph and let $\mathcal{H}$ and $\mathcal{K}$ be two families of connected subgraphs of $G$. An intersection hypergraph is the hypergraph $(\mathcal{H},\{\mathcal{H}_{K}\}_{K\in\mathcal{K}})$, i.e., a hypergraph on $\mathcal{H}$ such that each $K\in\mathcal{K}$ defines a hyperedge $\mathcal{H}_{K}$ consisting of the subgraphs in $\mathcal{H}$ intersecting a vertex of $K$. As in the dual setting, choosing a subset of $\mathcal{H}$ as terminals does not add any additional complication, and hence we define the intersection hypergraph on all of $\mathcal{H}$. Setting either $\mathcal{H}$ or $\mathcal{K}$ as singleton vertices yields, respectively, the primal and dual hypergraphs. An intersection support is a support for the intersection hypergraph.
@@
Our main result is the following: We show that if the host graph has genus $g$ and the subgraphs in $\mathcal{H}$ and $\mathcal{K}$ are cross-free, then there is an intersection support of genus $g$. This implies an equivalent result for the primal and dual hypergraphs. Figure 2 shows examples of primal, dual, and intersection supports.
@@
Figure 2: (a) and (b): Primal and Dual hypergraphs on the graph system $(G,\mathcal{H})$ where $G$ is the graph on vertices $\{a,b\ldots,f\}$, and $\mathcal{H}=\{H_{1},\ldots H_{4}\}$ with $H_{1}=\{a,b,c,d\}$, $H_{2}=\{c,d,e\}$, $H_{3}=\{a,b,f,e\}$, $H_{4}=\{a,b,c,e\}$; (c): Intersection hypergraph for $(G,\mathcal{H}^{\prime},\mathcal{K})$ with $\mathcal{H}^{\prime}=\{H^{\prime}_{1},\ldots,H^{\prime}_{4}\}$ and $\mathcal{K}=\{K_{1},\ldots,K_{3}\}$, where $H^{\prime}_{1}=\{a,b,c\}$, $H^{\prime}_{2}=\{a,b,c,e\}$, $H^{\prime}_{3}=\{c,e,f\}$, $H^{\prime}_{4}=\{b,e,d\}$, and $K_{1}=\{c,d\}$, $K_{2}=\{b,e,f\}$, $K_{3}=\{c,d,e\}$. Figures in (d)-(f) show the respective support graphs. Note: the pair of subgraphs $H^{\prime}_{3},H^{\prime}_{4}$ is not cross-free at vertex $e$ (see Definition 5).
@@ 1.2 Related Work
The notion of a planar analogue of a hypergraph was first suggested by Zykov [zykov], who defined a hypergraph to be planar if there is a plane graph on the vertices of the hypergraph such that for each hyperedge, there is a bounded face of the embedding containing only the elements of this hyperedge. Equivalently, a hypergraph is Zykov-planar iff its incidence bipartite graph is planar. Voloshina and Feinberg [voloshina1984planarity] introduced the notion of hypergraph planarity in the context of planarizing VLSI circuits (see the monograph by Feinberg et al., [feinberg2012vlsi] and references therein). The notion is now called a planar support. Johnson and Pollak [Johnson1987Hypergraph] showed that the problem of deciding if a hypergraph has a planar support is NP-hard.
@@
Since then, several authors have studied the question of deciding if a hypergraph admits a support from a restricted family of graphs. Tarjan and Yannakakis [tarjan1984simple] showed that we can decide in linear time if a hypergraph admits a tree support. Buchin et al., [buchin2011planar] showed polynomial time algorithms to decide if a hypergraph admits a support that is a path, a cycle, or a tree with bounded maximum degree. Further, the authors sharpen the result of Johnson and Pollak [Johnson1987Hypergraph] by showing that deciding if a hypergraph admits a support that is a 2-outerplanar [^4] A graph is 2-outerplanar if the graph can be embedded in the plane such that the vertices are on two concentric circles, and removing all vertices of the outer face results in an outerplanar graph. graph, is NP-hard. The problems of constructing a support with the fewest number of edges, or with minimum maximum degree have also been studied in [baldoni2007tera, onus2011minimum].
@@
As stated earlier, the existence of supports and related notions has been used in the analysis of algorithms under a local-search framework for packing and covering problems on geometric hypergraphs. Chan and Har-Peled [ChanH12] gave a PTAS for the Independent Set problem for an arrangement of pseudodisks in the plane. Mustafa and Ray [mustafa2010improved] gave a PTAS for the Hitting Set problem for a set of points and pseudodisks in the plane (both results apply more generally to k-admissible regions  [^5] A set of connected bounded regions $\mathcal{R}$ in the plane, each of whose boundary is a simple Jordan curve is $k$-admissible (for even $k$) if for any $^{\prime}\in\mathcal{R}, RRR\setminus R^{\prime}$ and $R^{\prime}\setminus R$ are connected, their boundaries are in general position, and intersect each other at most $k$ times. If $k=2$, the regions are called pseudodisks.[whitesides1990k] ). The results of [ChanH12, mustafa2010improved] were extended by Basu-Roy et al., [BasuRoy2018] who used the same framework to design a PTAS for the Set Cover and Dominating Set problems for the intersection graph of pseudodisks. Raman and Ray [RR18] showed the existence of a planar support for intersection hypergraphs of non-piercing regions (see Section 7 for the definition), and their result gave a unified analysis for all packing and covering problems described above. The local-search framework has also found applications in variants of the art gallery/terrain guarding problem [krohn2014guarding, DBLP:conf/wads/BandyapadhyayR17], where the geometric objects are defined implicitly as the visibility region of a guard, and the regions to be guarded are continuous regions.
@@
Cabello and Gajser [CabelloG14] showed that Independent Set and Vertex Cover admit a PTAS on graphs excluding a fixed minor, and Aschner et al., [DBLP:conf/walcom/AschnerKMY13] studied some packing and covering problems involving geometric non-planar graphs. Unlike their results, for the problems we study, we require additional work to show the existence of an appropriate local-search graph, which in some cases requires the construction of a support graph.
Besides the packing and covering problems, we also consider coloring problems on hypergraphs. Keller and Smorodinsky [KellerS18] considered intersection hypergraphs of two families of disks, and showed that such hypergraphs admit a conflict-free coloring [^6] In a conflict-free coloring, we want to color the vertices of the hypergraph such that each hyperedge has a uniquely colored vertex. with $O(\log n)$ colors. Keszegh [Keszegh20] generalized their result to intersection hypergraphs of two families of pseudodisks and showed that these hypergraphs admit a conflict-free coloring with $O(\log n)$ colors. The result holds for any hereditary family of hypergraphs if each hypergraph in the family admits a proper coloring using a constant number of colors [even_conflictfree]. Therefore, for certain geometric hypergraphs defined on a surface of constant genus, our result in Section 7.3 generalize the previous work of [even_conflictfree, KellerS18, Keszegh20, RR18].
@@ 2 Preliminaries
For a vertex $v\in V(G)$, we use $N_{G}(v)$ to denote the neighbors of $v$ in $G$ (or just $N(v)$ if $G$ is clear from context). We use $u\sim v$ or $e\sim v$ to denote respectively, if $u$ and $v$ are adjacent in $G$, or an edge $e$ incident to vertex $v$. For $S\subseteq V$, we use $G[S]$ to denote the subgraph of $G$ induced on $S$. We use $\mathcal{H}_{v}=\{H\in\mathcal{H}:v\in H\}$. Similarly, let $\mathcal{H}_{e}=\{H\in\mathcal{H}:e\in H\}$. We let $0pt(v)=|\{H\in\mathcal{H}:v\in H\}|$. Similarly, let $0pt(e)=|\{H\in\mathcal{H}:e\in H\}|$. We also use the notations $H\cap K$, $H\setminus K$, and $H\subseteq K$ to mean $V(H)\cap V(K)$, $V(H)\setminus V(K)$, and $V(H)\subseteq V(K)$ respectively for any two subgraphs $H,K$ of a graph $G$.
@@ 3.2 Algorithms for Packing and Covering Problems
For a graph system $(G,\mathcal{H})$, an intersection graph is the graph $D=(\mathcal{H},E)$, where $\{H,H^{\prime}\}\in E(D)$ if and only if $V(H)$ and $V(H^{\prime})$ intersect at a vertex of $G$.
@@ 4 Non-piercing and Cross-free Systems
Let $\mathcal{R}$ be an arrangement of connected and bounded regions in the plane, and $\partial(R)$ defines the boundary of $R\in\mathcal{R}$. Then $\mathcal{R}\setminus\bigcup_{R\in\mathcal{R}}\partial(R)$ splits the plane into a set $\mathcal{C}$ of connected components, each called a cell in the arrangement $\mathcal{R}$. Then the dual arrangement graph $G=(\mathcal{C},E)$ of $\mathcal{R}$ is a plane graph with vertex set $\mathcal{C}$ where two vertices are adjacent if and only if their corresponding cells are separated by boundary of some region in $\mathcal{R}$.
@@ 5 Vertex Bypassing
Let $(G,\mathcal{H})$ be a cross-free system with a cross-free embedding of $G$ with respect to $\mathcal{H}$. Suppose we apply VB(v)(v) to a vertex $v\in V(G)$. Then, each subgraph $H$ in $(G^{\prime},\mathcal{H}^{\prime})$ is connected. Further, VB(v)(v) can be done in time $O(|\mathcal{H}||V(G)|^{4})$.
@@
Consider an $x^{\prime}\in V(G^{\prime})$ and its corresponding vertex in $R_{G^{\prime}}(H^{\prime}_{1},H^{\prime}_{2})$. First, suppose $v\not\in V(H_{1})\cap V(H_{2})$. If $x^{\prime}\not\in\{u_{0},\ldots,u_{k-1}\}$, then the cyclic-pattern at $\tilde{x^{\prime}}$ in $R_{G^{\prime}}(H^{\prime}_{1},H^{\prime}_{2})$ is isomorphic to the cyclic-pattern at $\tilde{x}$ (replacing $H^{\prime}_{j}$ by $H_{j}$, $j=1,2$), where $x$ is the vertex in $G$ corresponding to $x^{\prime}$. Since $H_{1}$ and $H_{2}$ are cross-free at $x$, so are $H^{\prime}_{1}$ and $H^{\prime}_{2}$ at $x^{\prime}$. Since the subgraphs at $u_{i}$ are a subset of the subgraphs at $v$, it implies that no $u_{i}$ is contained in both $H_{1}$ and $H_{2}$. Hence, $H_{1}$ and $H_{2}$ are cross-free at each $u_{i}$.
Now, if $v\in V(H_{1})\cap V(H_{2})$, then for any vertex $y^{\prime}$ in $G^{\prime}$ so that its copy $y$ in $G$ is not connected to $v$ via a path that lies in $V(H_{1})\cap V(H_{2})$, the cyclic-pattern at $\tilde{y^{\prime}}$ and $\tilde{y}$ are isomorphic. If $y$ lies on a path in $V(H_{1})\cap V(H_{2})$ to $v$, then $\tilde{y}=\tilde{v}$ in $R_{G}(H_{1},H_{2})$. In this case, observe that the cyclic pattern at $\tilde{y^{\prime}}$ is a sub-sequence of the cyclic pattern at $\tilde{y}$ since $\tilde{y^{\prime}}=\tilde{u_{i}}$ for some $u_{i}\in H^{\prime}_{1}\cap H^{\prime}_{2}$. Since $H_{1}$ and $H_{2}$ are non-crossing at $v$, it implies that $H^{\prime}_{1}$ and $H^{\prime}_{2}$ are non-crossing at $y^{\prime}$.
Since $H^{\prime}_{1}$ and $H^{\prime}_{2}$ are non-crossing at all vertices of $G^{\prime}$, and $H_{1},H_{2}\in\mathcal{H}$ were arbitrary, it implies $(G^{\prime},\mathcal{H}^{\prime})$ is cross-free. ∎